# On diffusion-based generative models and their error bounds: The log-concave case with full convergence estimates

### v4

@@ C.2 Proof of the Main Result for the General Case
$\begin{split}&\leq\frac{\zeta}{2}\int_{0}^{t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\ \mathrm{d}s\\ &\quad+\frac{1}{2\zeta}\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}|b^{(k)}(T-s,\hat{\theta},\widehat{Y}_{s}^{\text{EM}})-b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ -\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\left(\widehat{Y}^{\text{EM},(j)}_{s}-\widehat{Y}^{\text{EM},(j)}_{\lfloor s/\gamma\rfloor\gamma}\right)|^{2}\ \mathrm{d}s\Bigg{]}\\ &\quad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\\ &\qquad\qquad\qquad\quad\times\left(\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}b^{(j)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor r/\gamma\rfloor\gamma})\ \mathrm{d}r\right)\mathrm{d}s\Bigg{]}\\ &\quad+\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right)\mathrm{d}s\right]\\ &\leq\frac{\zeta}{2}\int_{0}^{t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\ \mathrm{d}s+\frac{\mathsf{K}_{4}^{2}}{2\zeta}\int_{0}^{t}(1+2|T-s|^{\alpha})^{2}\ \mathbb{E}\left[|\widehat{Y}^{\text{EM}}_{s}-\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma}|^{4}\right]\ \mathrm{d}s\\ &\quad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\\ &\qquad\qquad\qquad\quad\times\left(\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}b^{(j)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor r/\gamma\rfloor\gamma})\ \mathrm{d}r\right)\ \mathrm{d}s\Bigg{]}\\ &\quad+\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}^{(j)}_{r}\right)\ \mathrm{d}s\right]\\ &\leq\frac{\zeta}{2}\int_{0}^{t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\ \mathrm{d}s+\gamma^{2}\frac{\mathsf{K}_{4}^{2}}{\zeta}t(1+4T^{2\alpha})C_{\mathsf{EMose},4}\\ &\quad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\\ &\qquad\qquad\qquad\quad\times\left(\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}b^{(j)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor r/\gamma\rfloor\gamma})\ \mathrm{d}r\right)\mathrm{d}s\Bigg{]}\\ &\quad+\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right)\mathrm{d}s\right].\end{split}$ (68)
@@
We proceed estimating the third and fourth term on the right-hand side of (68), separately.
@@
The third term is estimated using Young’s inequality with $\zeta\in(0,1)$, Lemma C.4, Remark 3.5, Remark 3.3 and Lemma C.1, and for any $t\in[0,T-\epsilon]$, we obtain that
@@
$\begin{split}&\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\\ &\qquad\qquad\qquad\times\left(\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}b^{(j)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor r/\gamma\rfloor\gamma})\mathrm{d}r\right)\mathrm{d}s\Bigg{]}\\ &\leq\frac{\zeta}{2}\int_{0}^{t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\ \mathrm{d}s\\ &\quad+\frac{2\gamma}{\zeta}M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\ \mathbb{E}\left[\int_{0}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sum_{j=1}^{M}|b^{(j)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor r/\gamma\rfloor\gamma})|^{2}\mathrm{d}r\ \mathrm{d}s\right]\\ &\leq\frac{\zeta}{2}\int_{0}^{t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\ \mathrm{d}s\\ &\quad+\frac{4\gamma}{\zeta}M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\qquad\qquad\times\int_{0}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\left[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))\mathbb{E}[|\widehat{Y}^{\text{EM}}_{\lfloor r/\gamma\rfloor\gamma}|^{2}]+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+\mathbb{E}[|\hat{\theta}|^{2}])\right]\mathrm{d}r\ \mathrm{d}s\\ &\leq\frac{\zeta}{2}\int_{0}^{t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\ \mathrm{d}s\\ &\quad+\frac{\gamma^{2}}{\zeta}4M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\times t\left[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))\sup_{0\leq s\leq t}\mathbb{E}[|\widehat{Y}^{\text{EM}}_{s}|^{2}]+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})\right]\\ &\leq\frac{\zeta}{2}\int_{0}^{t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\ \mathrm{d}s\\ &\quad+t\frac{\gamma^{2}}{\zeta}4M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\times\left[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})\right].\end{split}$ (69)
@@
We proceed with the estimate of the fourth term on the right-hand side of (68). For $k=1,\dots,M$, we note
@@
By using (70), Cauchy–Schwarz inequality, Young’s inequality, Assumption 3 and Lemma C.4, we have
@@
$\begin{split}&\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}^{(j)}_{r}\right)\mathrm{d}s\right]\\ &=\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\left(Y_{\lfloor s/\gamma\rfloor\gamma}^{\text{aux},(k)}-\widehat{Y}_{\lfloor s/\gamma\rfloor\gamma}^{\text{EM},(k)}\right)\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}^{(j)}_{r}\right)\mathrm{d}s\right]\\ &\qquad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(\int_{\lfloor s/\gamma\rfloor\gamma}^{s}(b^{(k)}(T-r,\hat{\theta},Y_{r}^{\text{aux}})-b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{r}^{\text{EM}}))\ \mathrm{d}r\right)\\ &\qquad\qquad\qquad\qquad\times\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}^{(j)}_{r}\right)\ \mathrm{d}s\Bigg{]}\\ &\qquad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(\int_{\lfloor s/\gamma\rfloor\gamma}^{s}(b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{r}^{\text{EM}})-b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}}))\ \mathrm{d}r\right)\\ &\qquad\qquad\qquad\qquad\times\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}^{(j)}_{r}\right)\ \mathrm{d}s\Bigg{]}\\ &\qquad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(\int_{\lfloor s/\gamma\rfloor\gamma}^{s}2(s^{(k)}(T-r,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}})-s^{(k)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}}))\ \mathrm{d}r\right)\\ &\qquad\qquad\qquad\qquad\times\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}^{(j)}_{r}\right)\ \mathrm{d}s\Bigg{]}\\ &\leq\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\gamma^{-1/2}|b^{(k)}(T-r,\hat{\theta},Y_{r}^{\text{aux}})-b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{r}^{\text{EM}})|\mathrm{d}r\\ &\qquad\qquad\qquad\qquad\times\gamma^{1/2}\left|\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right|\mathrm{d}s\Bigg{]}\\ &\qquad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}|b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{r}^{\text{EM}})-b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}})|\mathrm{d}r\\ &\qquad\qquad\qquad\quad\times\left|\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right|\mathrm{d}s\Bigg{]}\\ &\qquad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}2|s^{(k)}(T-r,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}})-s^{(k)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}})|\mathrm{d}r\\ &\qquad\qquad\qquad\quad\times\left|\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right|\mathrm{d}s\Bigg{]}\end{split}$

### v5

@@ D.2 Proof of the Main Result for the General Case
$\begin{split}&\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ -\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\left(\widehat{Y}^{\text{EM},(j)}_{s}-\widehat{Y}^{\text{EM},(j)}_{\lfloor s/\gamma\rfloor\gamma}\right)|^{2}\ \mathrm{d}s\Bigg{]}\\ &\quad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\\ &\qquad\qquad\qquad\quad\times\left(\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}b^{(j)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor r/\gamma\rfloor\gamma})\ \mathrm{d}r\right)\mathrm{d}s\Bigg{]}\\ &\quad+\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right)\mathrm{d}s\right]\\ &\leq\frac{\zeta}{2}\int_{0}^{t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\ \mathrm{d}s+\frac{\mathsf{K}_{4}^{2}}{2\zeta}\int_{0}^{t}(1+2|T-s|^{\alpha})^{2}\ \mathbb{E}\left[|\widehat{Y}^{\text{EM}}_{s}-\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma}|^{4}\right]\ \mathrm{d}s\\ &\quad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\\ &\qquad\qquad\qquad\quad\times\left(\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}b^{(j)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor r/\gamma\rfloor\gamma})\ \mathrm{d}r\right)\ \mathrm{d}s\Bigg{]}\\ &\quad+\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}^{(j)}_{r}\right)\ \mathrm{d}s\right]\\ &\leq\frac{\zeta}{2}\int_{0}^{t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\ \mathrm{d}s+\gamma^{2}\frac{\mathsf{K}_{4}^{2}}{\zeta}t(1+4T^{2\alpha})C_{\mathsf{EMose},4}\\ &\quad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\\ &\qquad\qquad\qquad\quad\times\left(\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}b^{(j)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor r/\gamma\rfloor\gamma})\ \mathrm{d}r\right)\mathrm{d}s\Bigg{]}\\ &\quad+\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right)\mathrm{d}s\right].\end{split}$ (73)
@@
We proceed estimating the third and fourth term on the right-hand side of 73, separately.
@@
The third term is estimated using Young’s inequality with $\zeta\in(0,1)$, Lemma 23, Remark 5, Remark 3 and Lemma 20, and for any $t\in[0,T-\epsilon]$, we obtain that
@@
$\begin{split}&\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\\ &\qquad\qquad\qquad\times\left(\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}b^{(j)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor r/\gamma\rfloor\gamma})\mathrm{d}r\right)\mathrm{d}s\Bigg{]}\\ &\leq\frac{\zeta}{2}\int_{0}^{t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\ \mathrm{d}s\\ &\quad+\frac{2\gamma}{\zeta}M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\ \mathbb{E}\left[\int_{0}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sum_{j=1}^{M}|b^{(j)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor r/\gamma\rfloor\gamma})|^{2}\mathrm{d}r\ \mathrm{d}s\right]\end{split}$
$\begin{split}&\leq\frac{\zeta}{2}\int_{0}^{t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\ \mathrm{d}s\\ &\quad+\frac{4\gamma}{\zeta}M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\qquad\qquad\times\int_{0}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\left[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))\mathbb{E}[|\widehat{Y}^{\text{EM}}_{\lfloor r/\gamma\rfloor\gamma}|^{2}]+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+\mathbb{E}[|\hat{\theta}|^{2}])\right]\mathrm{d}r\ \mathrm{d}s\\ &\leq\frac{\zeta}{2}\int_{0}^{t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\ \mathrm{d}s\\ &\quad+\frac{\gamma^{2}}{\zeta}4M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\times t\left[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))\sup_{0\leq s\leq t}\mathbb{E}[|\widehat{Y}^{\text{EM}}_{s}|^{2}]+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})\right]\\ &\leq\frac{\zeta}{2}\int_{0}^{t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\ \mathrm{d}s\\ &\quad+t\frac{\gamma^{2}}{\zeta}4M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\times\left[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})\right].\end{split}$ (74)
@@
We proceed with the estimate of the fourth term on the right-hand side of 73. For $k=1,\dots,M$, we note
@@
By using 75, Cauchy–Schwarz inequality, Young’s inequality, Assumption 3.b and Lemma 23, we have
@@
$\begin{split}&\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}^{(j)}_{r}\right)\mathrm{d}s\right]\\ &=\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\left(Y_{\lfloor s/\gamma\rfloor\gamma}^{\text{aux},(k)}-\widehat{Y}_{\lfloor s/\gamma\rfloor\gamma}^{\text{EM},(k)}\right)\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}^{(j)}_{r}\right)\mathrm{d}s\right]\\ &\qquad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(\int_{\lfloor s/\gamma\rfloor\gamma}^{s}(b^{(k)}(T-r,\hat{\theta},Y_{r}^{\text{aux}})-b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{r}^{\text{EM}}))\ \mathrm{d}r\right)\\ &\qquad\qquad\qquad\qquad\times\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}^{(j)}_{r}\right)\ \mathrm{d}s\Bigg{]}\end{split}$